Sol Summary – November 18th

Sol: 5
Summary Title: Misstep on Mars
Author’s name: Jas Purewal
Mission Status: Nominal
Sol Activity Summary: Crew headed out for a morning EVA to the junction of Cow Dung Rd and Brahe Hwy to test the recently serviced rovers. Opportunity overheated on the return leg of EVA. The crew waited for ~ 10 mins but Oppy still didn’t start. They left Oppy in the field and returned back to the hab. As there was more time left in the EVA, the EVA crew learnt that they should have waited another 20 mins for Oppy to cool before returning. Lesson Learned. The rover keys were brought back to the hab and left in the RAM for the outpost team to retrieve.
In the afternoon the crew reviewed the EVA and lessons learned. Discussed mission and EVA planning. Did VR first aid training and practised splinting a broken ankle. We also spent some much needed relaxation time, played chess and made Mac n Cheese (with a kick). The crew Biologist made LB and PDA plates in the Science Dome.
Look Ahead Plan: One morning EVA planned from 8am to 10:30am. XO, HSO and Biologist taking Spirit and Curiosity to explore “The Peanut” via Cow Dung and Galileo Rd to collect biological samples with a focus on lichen and any soil around plants. In the afternoon we will do more VR first aid training and team development training. The crew biologist will analyse the collected samples and use the LB plates to grow collected samples in the Science Dome.
Anomalies in work: None
Weather: Rather nice. 42F, sunny, clear. No rain.
Crew Physical Status: Good
EVA: EVA 6 – Crew headed out for a morning EVA at 9:30am to the junction of Cow Dung Rd and Brahe Hwy to test the recently serviced rovers. Opportunity overheated on the return leg of the EVA. The crew waited for ~ 10 mins but Oppy still didn’t start. They left Oppy in the field and returned back to the hab. As there was more time left in the EVA, the EVA crew learnt that they should have waited another 20 mins for Oppy to cool before returning. Lesson Learned. The rover keys were brought back to the hab and left in the RAM for the outpost team to retrieve.
Reports to be filed:
Commander’s report
EVA report
EVA request
Operations report
Sol Summary report
Journalist report

Support Requested: Retrieval for Opportunity rover

Operations Report – November 18th

SOL: 5
Name of person filing report: Jas Purewal
Non-nominal systems: Opportunity overheated.
Notes on non-nominal systems: NA
Spirit rover used: Yes
Hours: (before EVA) 198.2
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging) 56%
Currently charging: Yes
Opportunity rover used: Yes
Hours: 103.7
Beginning charge: 100%
Ending charge: 77%
Currently charging: No
Curiosity rover used: Yes
Hours: 210.4
Beginning charge: 100%
Ending charge: 54%
Currently charging: Yes
Perseverance rover used: No
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: NA
General notes on rovers: Opportunity overheated. He was abandoned in the field to be rescued at a later time.
Summary of Hab operations: put notes here
WATER USE: 50 gallons
Water (static tank): 350 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off) off
Toilet tank emptied: Yes
Summary of internet: Good
Summary of suits and radios: optimal
Summary of GreenHab operations: None
WATER USE: 5.02 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: None
Dual split: (Heat or AC, On or Off) Heat, Off
Summary of RAM operations: (Tools used, work done) Plumbing adhesive and tape currently in main hab. We put Opportunity’s keys in the RAM to be retrieved.
Summary of any observatory issues: NA
Summary of health and safety issues: none
Questions, concerns and requests to Mission Support: Still waiting on more trash bags please for toilet bins and for general bins. Pillows please.

Journalist Report – November 18th

Sol: 5
Title: Mars hits back
Author: Izabela Shopova, Crew Journalist
Another gold and blood sunrise. Mars is trying so hard to win our hearts over. The alien landscape with its breathtaking beauty is slowly getting under our skin, just as the martian dust is making itself at home in our nostrils, ears and hair.

The crew slept well for the whole night – we are finally getting over the jetlag and the adrenalin rush of the first days, and adjusting to our martian home away from home.

Today is the 8th day since we all left our respective homes and hit the road. We have been away from family and loved ones for a full week now and we are starting to feel it. Last night the conversations revolved around our personal life. Photos of partners, family and pets were shared, and funny stories were told. Mars is starting to feel real now.

We appear to have established some routines of our own. Like a family, we always gather around the table for breakfast, even if not everyone eats. SHO checks how everyone feels, we chat for a bit and then we go over our plans for the day, confirming everyone’s responsibilities, who is cooking, who goes on EVA and what other projects we need to work on. Dinner is another time when we get together and once again enjoy a shared meal and a lively conversation, as we discuss the lessons learned and our plans for the next day.

Our morning EVA was going really well today, we felt calm and confident, following our procedures, ticking EVA objectives on the way and generally enjoying the feeling that our crew works like a well-oiled machine when suddenly a disaster struck. Mars hit back. “Opportunity” overheated and stopped, flashing an angry, bright red light on its dashboard.

The incident happened while we were out of radio contact with the Hub, so after two unsuccessful attempts to restart the engine we decided to abandon the rover on the road and return to the Hab early, prioritizing the crew’s health and safety over the vehicle recovery.

“Opportunity was rescued by the outpost team a bit later and upon reconsidering our options and possible responses to the emergency situation, we all agreed that abandoning the rover while we still had time to make additional rescue attempts was probably an error. And potentially a costly one on Mars. The Red Planet taught us another valuable lesson. The crew spent the afternoon reconstructing and studying the details of the EVA, considering the different options and the lessons learned, as well as planning objectives for the rest of the mission. Out here, on the Red Planet, we learn something new and valuable everyday. Mars tests and challenges us, helping us to learn and grow.

Today was no exception.

GreenHab Report – November 18th

Sol 5
GreenHab Officer: Caitlyn Hubric
Environmental control: (heater, fan, door open?) heater on, door closed, 20% humidity
Average temperatures: 43.3ºC/13.5ºC
Hours of supplemental light: N/A
Daily water usage for crops: 18 L ≈ 4.76 gal
Daily water usage for research and/or other purposes: 1 L ≈ 0.26 gal
Water in Blue Tank ~277.3 gallons
Time(s) of watering for crops: twice, morning and night, 9am-5pm
Changes to crops:
-Beans and sugar snap peas look very healthy!
-The dill in the blue bin is starting to look very full and is growing well.
-Might be able to harvest the dill and the other assorted kale within the next week.
-My decomposition research is thriving! I already have mycelium on the top surface of my substrate. I am waiting for a stronger presence before I encourage fruiting.
Harvest: (include which crop and mass in grams) N/A
Support/supplies needed: N/A

EVA Report – November 18th

EVA # 6
Author: Izabela Shopova (Narrative), Elizabeth Balga (Hotwash Notes)
Purpose of EVA: Rover work (per Dr. Rupert). Engineering sample (rock) collection.
Start time: 9:30 AM
End time: 11:00 AM
The EVA started on time and as planned the crew was driving 3 rovers – ‘Opportunity’, ‘Spirit’ and ‘Curiosity’ – to test them after their servicing.
Weather was clear and sunny, and the rovers performed well. Crew noted the position where radio connection with the Hub was lost and switched to channel 2. Reaching the established goal of the expedition – the intersection of Cow Dung Road and Brahe Rd – the crew stopped, noted the battery performance, took photos and explored the area for 15 minutes.
On the return back to the Hub the crew stopped after crossing a river bed to explore and take photos of local vegetation for the crew biologist. Soon after continuing with the return drive ‘Opportunity’ lost power and stopped. Red light came up on the dashboard indicating overheating.
Hotwash Notes:
Bottom line up front — In the future, the team should better utilize available time on EVA in the event of an anomaly. Crew was prioritizing safety and sim, but should also prioritize mission critical equipment based on available time.
EVA team initiated return from Brahe intersection around 10:02 AM
Oppy’s battery was around ~75% upon initiating return, best battery performance of the 3 rovers
EVA team stopped at mud flat on the return trip. Departed mud flat around 10:12 AM
En route back to habitat, Oppy broke down around ~10:15 AM
Oppy was located at the rear of convoy and broke down on a slope (slightly downward)
Driver noted that the red light on the dashboard was illuminated. May have turned on earlier but difficult to see with spacesuit, glare, and steering wheel
Driver waited with rover for ~10 minutes TOTAL while crewmates returned to spot & pursued troubleshooting
Tried to start rover again two times (about 5 mins between each attempt), no joy
There were no communications with the hab or outpost at the location of rover breakdown. The
Team wanted to re-established comms in case of additional instructions from HABCOM or Outpost, but did not want to split up and strand a crew member per protocol.
Ultimately team had additional time (20 mins) to spare before EVA end, so could have spent another 10 minutes waiting before abandoning
Note: Times validated during hotwash with HABCOM and photos
Additionally, had tow rope but no training/protocol for towing
Also concern with towing for risk to remaining rovers since breakdown occurred in gulley
Crew left tow rope with abandoned rover to assist with recovery
Crew brought keys back to station per protocol
Crew abandoned the rover and initiated return around 10:25 AM
Crew entered the airlock around ~10:35 AM
Destination: Intersection of Cow Dung Rd and Brahe Hwy.
Coordinates (use UTM WGS 84): 518000, 4255000
Participants: CDR, COMM, ENGR
Road(s) and routes per MDRS Map:
North on Cow Dung Rd (0110) towards Brahe Hwy on Rover
Mode of travel: Rover, Walking (Oppy, Curi, Spirit)

Commander Report – November 18th

Sol: 5
Summary Title: Misstep on Mars
Author’s name: Dr. Jenni Hesterman, Commander
Mission Status: Nominal
Commander Report:

The crew was up early and enjoyed a light breakfast while discussing the day ahead.

In the only scheduled EVA for the day, 3 crew members took 3 rovers out to test them after recent servicing. They proceeded up Cow Dung Road to Brahe Highway, the turnaround point. On the return trip to the station, one of the rovers overheated and stopped working in a gulley. The other crew members safely parked and made their way on foot down to the location. After a few attempts to restart the rover, the crew decided to leave it behind and return to base 25 minutes before the end of the mission to assess the situation. As the rover is a precious resource, the crew should have spent the extra time waiting to see if the vehicle would cool off. We continue to learn each day and know these mistakes and experiences will inform our future work.

Over a lunch of baked macaroni and cheese, the crew discussed lessons learned and updated the EVA procedures checklists. They enjoyed a quiet afternoon in the station, reading, playing space chess and interacting with Paro, an artificial intelligence therapeutic robot in the form of a baby seal.

We’re excited to get back out on an EVA tomorrow morning! We will accomplish first aid training and team building activities in the afternoon

Copyright © The Mars Society. All rights reserved. | Main Site